OpenStudy (anonymous):

Find all roots of 4x^2-7x+3

OpenStudy (anonymous):

\[\frac{ -b^2\pm \sqrt{b^2-4ac} }{ 2a }\] where \( ax^2+bx+c=0 \)

OpenStudy (anonymous):

I come up with (4x+3)(1x-1) 4x-3=0, x=3/4 3x-3=0 x =-1 Doe those answers look correct?

OpenStudy (anonymous):

no, i assume you mixed up a sign somewhere

OpenStudy (anonymous):

Last one should have been x-1=0 so x=1